A car moving at 40 km/h is to be stopped applying brakes in the next 4.0 m. If the car weighs 2000 kg what average force must be applied on it?
please...

Grade:11

1 Answers

jakira
7 years ago
1 km = 1000 m, 1 h = 3600 s, 1 km/h = 1000/3600 = 5/18 m/s 
v = 40 * 5/18 = 100/9 
This is approximately 11.1 m/s. 
vf^2 = vi^2 + 2 * a * d, vf = 0, d = 4 
0 = (100/9)^2 + 2 * a * 4 
a = -(100/9)^2 ÷ 8 
This is approximately -15.432 m/s^2
 
Average force = mass * acceleration = 2000 * -(100/9)^2 ÷ 8 
This is approximately -30,864 N.